OpenStudy (anonymous):

Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y A. The solution is (1, 5) B. The solution is (21, 0) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

i am thinking a or c, more leaning to c though plz help I WILL GIVE MEDAL AND FAN!

OpenStudy (anonymous):

y + 2x = 7 14 – 4x = 2y Simplify: 2y = 14 - 4x y = 7 - 2x Substitute: y + 2x = 7 (7 - 2x) + 2x = 7 7 = 7 As this shows any value of x and y are irrelevant... C. There are an infinite number of solutions.

OpenStudy (anonymous):

omg thank you so much! mind helping me with more? i will tag u

OpenStudy (anonymous):

sure :)